# Redshift leverage for the search of GRB neutrinos affected by quantum properties of spacetime

### v1

@@ I Introduction and new selection criteria
The current generation of neutrino telescopes was devised to mark the start of neutrino astrophysics, and indeed IceCube firmly established the observation of cosmological neutrinos [1], but so far not much astrophysics has been done with neutrinos: we only have a single case of observation of a source in neutrinos, the case of the active galaxy NGC 1068 for which IceCube reported robust statistical evidence of neutrino observations [2]. With respect to pre-IceCube expectations what is clearly missing are observations of neutrinos from GRBs (gamma-ray bursts): the prediction of a neutrino emission associated with GRBs is generic within the most widely accepted astrophysical models [3] and pre-IceCube studies estimated about a handful of GRB-neutrino observations per year of operation [4, 5, 6, 7], but as IceCube nears 15 years of operation still not a single GRB neutrino has been observed. The simplest (and most likely) explanation of this huge underperformance of IceCube is that our GRB models must be revised in such a way to produce a much lower neutrino flux, but it is intriguing that in principle our failure to observe GRB neutrinos could also be attributed to some plausible quantum properties of spacetime which had already been investigated for independent reasons in the quantum-gravity literature: IceCube searches might have failed to find GRB neutrinos because they assume that GRB neutrinos should be detected in very close temporal coincidence with the associated gamma rays, but a sizable mismatch between GRB-neutrino detection time and the time of detection of the electromagnetic GRB signal would be expected in presence of “in-vacuo dispersion”, a much-studied effect such that quantum properties of spacetime slow down [^1] While the theoretical prejudice of most quantum-spacetime researchers favours effects that slow down particles, there is no conclusive theoretical argument ruling out that quantum-spacetime effects might instead speed up particles. We offer some data-based observations relevant for this issue in parts of Sec. V. particles proportionally to their energies (see, e.g., Refs. [8, 9, 10, 11, 12, 13, 14, 15, 16, 17] and references therein).
@@
where $\Delta t$ is the contribution to the travel time of the neutrino from quantum-gravity-induced in-vacuo dispersion, $M_{P}$ denotes the Planck scale ($\sim 10^{28}eV$), and $D(z)$ is a function of the redshift $z$ of the GRB associated to the neutrino,
@@
We shall mainly focus on the scenario with pure systematic effects, in which $\eta$ is simply a parameter to be determined experimentally, but in a few points we shall offer remarks relevant for the more general case of a mixture of systematic and fuzzy effects [16, 17, 24], in which $\eta$ is described as $\eta_{0}+\delta\eta$, where $\eta_{0}$ is a fixed parameter (taking the same value in all neutrino observations) while $\delta\eta$ is an additional contribution that differs from one instance to another of neutrino observation and vanishes on average.
@@
As observed in some previous studies [18, 19, 20, 21, 22, 25, 26] the search of GRB neutrinos affected by in-vacuo dispersion must necessarily rely on a statistical approach: we could never be sure that a certain neutrino is a GRB neutrino, but there is a chance to have at some point a collection of GRB-neutrino candidates (neutrinos that could be associated with a GRB, assuming in-vacuo dispersion) large enough to be sure that at least some of those neutrinos actually are GRB neutrinos, affected by in-vacuo dispersion. This is due to the size of the time window that one must adopt in order to investigate the hypothesis of in-vacuo dispersion. Standard searches of GRB neutrinos (assuming $\eta=0$, no in-vacuo dispersion) can assume that the neutrino and the electromagnetic signal from the GRB travel essentially at the same speed, and therefore their GRB-neutrino candidates are looked for within a tiny time window, rendering background issues insignificant. For $\eta\neq 0$, also taking into account that the neutrino energy will have some uncertainty (and however one will inevitably always end up testing an hypothesis of $\eta$ taking values in a certain range), the difference in observation time between the neutrino and the electromagnetic signal would be sizeably uncertain, and the time window for searches of GRB-neutrino candidates would have to be correspondingly large, resulting in background issues such that one could not be sure that a specific GRB-neutrino association is correct.
@@
Evidently the effectiveness of such a statistical approach depends rather crucially on the criteria used for the selection of candidate GRB neutrinos. And our main objective here is to test a novel proposal for these criteria. Testing alternative strategies of analysis on presently-available IceCube data also acquires additional importance because it can set the stage for a later, more mature, phase of the research program, eventually using also data from KM3Net [27] and IceCube-Gen2  [28].
@@
Previous searches [18, 19, 20, 21, 22, 25, 26] of GRB neutrinos looked for neutrinos within a time window of fixed size (neutrino-energy-independent) after the GRB onset, and put on the same footing both GRBs whose redshift has not been measured (then having to estimate that redshift crudely, since the conjectured effect is redshift dependent) and GRBs of known redshift. Including GRBs whose redshift has not been measured has the advantage of a larger number of candidate GRB neutrinos, but of course renders the analysis vulnerable to the assumptions made to roughly estimate the redshifts. Moreover, the fixed-size time window, while easily handled computationally, imposes a restriction of the analysis to a corresponding limited range of neutrino energies: since the sought effect grows linearly with energy, a time window adapted to energies of, say, 100 TeV will inevitably be too small for neutrinos with energy much greater than 100 TeV, and in an appropriate sense it would also be too large a time window for neutrinos of energy much smaller than 100 TeV (when the time window is much wider than the one really needed by the sought effect, the analysis ends up being dominated by background). In spite of these limitations, Ref. [26], the latest such study (of which some of us were authors), provided the estimate [^2] As it will be clearer for our readers when, later in this manuscript, we derive an analogous $p$-value for our novel selection criteria, this $p$-value, which in [26] was called false alarm probability, is the probability that all the found GRB-neutrino candidates actually are background neutrinos (probability of no signal), while $\eta=21.7\pm 4.5$ is the estimate of $\eta$ obtained from the found GRB-neutrino candidates. $\eta=21.7\pm 4.5$ with a $p$-value $P_{[26]}=0.007$.
@@
We here take the results of Ref. [26] as a starting point for testing a novel strategy of analysis. Rather than adopting a fixed-size time window and correspondingly restricting the neutrino-energy range, we consider all the neutrinos in the sample, regardless of the energy, but restrict the GRB catalogue to GRBs of known redshift. Any GRB and neutrino times of arrival are then regarded as compatible if they satisfy Eq. (1) for some $\eta\in\left[12.7,30.7\right]$ (the two-standard-deviation interval obtained from the estimate $\eta=21.7\pm 4.5$ found in Ref. [26]) within two standard uncertainties in the neutrino visible energy. As here shown in Sec. IV, the fact that, by relying exclusively on GRBs of known redshift, we find fewer GRB-neutrino candidates is more than compensated, for what concerns the overall statistical significance, by the sharper setup of the analysis which is available when the redshift of the GRBs is known.
@@ II IceCube neutrinos and GRBs of known redshift
Our analysis relies on the latest HESE data release by IceCube (7 September 2023), available at [31]. As done in previous analogous studies [18, 19, 20, 21, 22, 26], we restrict to shower events, since track events have very poor energy estimates (and of course the quality of the energy information is crucial for in-vacuo-dispersion studies) [40, 42].
A sizable part of the effort we devoted to this study was aimed at compiling a reliable catalogue of GRBs with actually measured (and not just estimated or loosely constrained) redshifts. In fact, we found that all already available GRB catalogues either imposed further constraints on the GRB properties (narrower observation window, restriction to specific GRB observatories, GRB type, etc.) or were not sufficiently accurate (misreporting or omitting direction or redshift data, not taking into account the latest relevant GCNs, etc.). The resulting list is reported here in Appendix A. In about 90% of cases it agrees with an analogous general-purpose catalogue, not restricted to known-redshift GRBs, mantained by the IceCube collaboration [32]. The previous study [26] relied on this resource, but, after checking one-by-one each of its known-redshift entries, we noticed that some of the information relevant for our analysis was inaccurately reported, probably due to the automated population of the database. In most cases this was quickly corrected according to the web catalogue [33] mantained by Jochen Greiner at the MPE, which we found to be the most accurate and comprehensive among publicly available lists. Checking by hand all the relevant GCNs and consulting additional catalogues of known-redshift GRBs, we were then able to correct the few errors left and also identify a few known-redshift GRBs whose redshift is not reported in Greiner’s table (see Appendix A for details).
@@ III Simulated data for the statistical-significance analysis and background estimate
As done in the previous neutrino studies [^3] The interested reader can find applications of this method of statistical analysis to studies not involving neutrinos, e.g., in Refs. [30, 34, 35, 36, 37, 38]. of Refs. [18, 19, 20], we do this by producing simulated data through transformations of the real data which reliably preserve their morphology while washing away any possible correlations of the type predicted by Eq.(1) between the times of arrival, the neutrino energy, and $D(z)$.
@@
$\bullet$ the directions of known-redshift GRBs are independent of galactic longitude, but are correlated to some extent with galactic latitude, as the dust lying around the galactic plane makes it much more difficult to determine the redshift of a GRB in that region; also, we do not act on the GRB observation times because we find that the detection rate of known-redshift GRB is declining with time, probably another selection effect due to older events having been more likely followed by the long-term, host-galaxy observations which are often needed to determine the redshift of a GRB.
As a first application of our simulated data we can estimate the background that is expected with the criteria adopted by our strategy of analysis. For this purpose, we generated $10^{5}$ instances of simulated data and we used them to estimate the expected number of background neutrinos that would accidentally be associated with a GRB according to the time-energy and direction compatibility conditions specified above ($\eta\in[12.7,30.7]$ and $\mathcal{S}_{cut}=1/4\pi$). We find that, on average, our criteria should pick up 1.02 background neutrinos.
@@ IV Main analysis
Equipped with the preparatory work reported in the previous sections, we are now ready to discuss our main analysis. We begin by looking for GRB-neutrino candidates, using the data described in Sec. II. For each GRB-neutrino pair selected by our directional criteria ($\mathcal{S}>1/4\pi$) we describe the $\Delta t$ of Eq.(1) as the difference between the neutrino observation time [^4] If there was no in-vacuo dispersion, a GRB neutrino would be observed (nearly-)simultaneously with the GRB photons. We attribute the whole of the time-of-arrival difference to the $\Delta t$ of the neutrino, since photons observed from GRBs are of much lower energies than our neutrinos and the effect we are studying depends linearly on energy. Some quantum-spacetime scenarios predict in-vacuo dispersion with different magnitude for photons and neutrinos (and if the magnitude was much stronger for photons this assumption of our analysis would not be satisfied); it is however noteworthy that effective-field-theory approaches to quantum properties of spacetime predict that the type of pure in-vacuo dispersion here studied is actually forbidden for photons whereas it is allowed for neutrinos [15, 39]. and the observation time of the GRB, which we label as $\Delta t_{candidate}$, and then we use Eq.(1) and the known redshift of the GRB, which we label as $z_{GRB}$, to convert the requirement that $\eta$ should be within the interval $[12.7,30.7]$ into a corresponding range of allowed values for the neutrino energy. We then consider the GRB-neutrino pair a “GRB-neutrino candidate” if the energy of the neutrino is compatible with that energy range within two standard deviations (assuming 10% uncertainty in the energy of the neutrino [40]). We find that there are four such GRB-neutrino candidates. Fig. 1 provides a visual characterization of the fact that the properties of these four GRB-neutrino candidates match rather well the expectations of the in-vacuo-dispersion scenario of Eq. (1).
@@
The fact that our findings reported in Fig. 1 fit very well with the estimate $\eta=21.7\pm 4.5$, which resulted from the analysis of Ref. [26], is noteworthy since, because of the different criteria adopted for the selection of GRB-neutrino candidates, the two analyses only share one GRB-neutrino candidate, the one shown as a violet square in Fig. 1. The analysis leading to the estimate $\eta=21.7\pm 4.5$, in Ref. [26] ended up focusing on seven GRB-neutrino candidates, but only one of them involved a GRB whose redshift has been measured.

### v2

@@ I Introduction and new selection criteria
The current generation of neutrino telescopes was devised to mark the start of neutrino astrophysics, and indeed IceCube firmly established the observation of cosmological neutrinos [1], but so far not much astrophysics has been done with neutrinos: there is only evidence of a neutrino signal from the active galaxy NGC 1068 [2] and from the flaring blazar TXS 0506+056 (see, e.g., Refs. [3, 4]). With respect to pre-IceCube expectations what is clearly missing are observations of neutrinos from GRBs (gamma-ray bursts). The prediction of a neutrino emission associated with GRBs is generic within the most widely accepted astrophysical models [5] and pre-IceCube studies estimated about a handful of GRB-neutrino observations per year of operation [6, 7, 8, 9]; yet, as IceCube nears 15 years of operation, still not a single GRB neutrino has been reported. The simplest (and most likely) explanation of this huge underperformance of IceCube is that our GRB models must be revised in such a way to produce a much lower neutrino flux, but it is intriguing that in principle our failure to observe GRB neutrinos could also be attributed to some plausible quantum properties of spacetime which had already been investigated for independent reasons in the quantum-gravity literature: IceCube searches might have failed to find GRB neutrinos because they assume that a GRB neutrino should be detected in very close temporal coincidence with the associated gamma rays, but a sizable mismatch between the neutrino detection time and the time of detection of the electromagnetic GRB signal would be expected in presence of “in-vacuo dispersion”, a much-studied effect such that quantum properties of spacetime slow down [^1] While the theoretical prejudice of most quantum-spacetime researchers favors effects that slow down particles, there is no conclusive theoretical argument ruling out that quantum-spacetime effects might instead speed up particles. We offer some data-based observations relevant for this issue in parts of Sec. V. particles proportionally to their energies (see, e.g., Refs. [10, 11, 12, 13, 14, 15, 16, 17, 18, 19] and references therein).
@@
where $\Delta t$ is the contribution to the travel time of the neutrino due to quantum-gravity-induced in-vacuo dispersion, $E$ is the neutrino energy, $M_{P}$ denotes the Planck scale ($\sim 10^{16}$ TeV), and $D(z)$ is a function of the redshift $z$ of the source,
@@
We restrict to the scenario with purely systematic effects, in which $\eta$ is simply a parameter to be determined experimentally. When a larger collection of astrophysical neutrinos becomes available (and therefore analyses such as ours can attempt multi-parameter investigations), it will be possible to consider the more general case of a mixture of systematic and fuzzy effects [18, 19, 26], in which $\eta$ is described as a random variable (taking different values for different neutrinos) with mean $\eta_{0}$ and dispersion $\delta\eta$. We also postpone the extension of our phenomenological model to allow for the possibility that there might be intrinsic at-the-source effects causing a sizable offset between the typical neutrino emission time and the GRB onset. A first step for modeling the impact of such intrinsic effects on in-vacuo dispersion analyses would be to add to Eq.(1) an extra term of the form ${\bar{\tau}}(1+z)$, where $\bar{\tau}$ is the average neutrino emission time offset with respect to the GRB onset and the factor $(1+z)$ accounts for ordinary time dilation, but investigations of such a refined model need more astrophysical neutrinos than presently available.
@@
Indeed, in-vacuo dispersion for photons has been investigated very intensely over the last 15 years (see, e.g., Refs. [18, 19] and references therein), with some studies even reaching sensitivity of the order of the Planck scale (see, e.g., Ref. [27]). It is difficult to say exactly which values of $\eta$ have been excluded for photons, since the magnitude of in-vacuo-dispersion effects could be comparable to known (but poorly understood) intrinsic spectral lags and a delayed emission of the relevant photons cannot be ruled out (some of the most energetic photons from a GRB could be emitted within the long afterglow phase). Still, a balanced perspective is that, for photons, we are close to Planck-scale sensitivity [18, 19].
The investigation of in-vacuo dispersion for neutrinos has been lagging far behind: the current bound on in-vacuo dispersion for neutrinos is still based on the historic observation of neutrinos from the SN1987a supernova, and only amounts to $\eta\lesssim 10^{9}$ [28]. This is very unsatisfactory, especially in light of the fact that the quantum-spacetime literature provides motivation for studying in-vacuo dispersion for photons and neutrinos separately. In particular, quantum-spacetime models based on Planck-scale discretization predict in-vacuo dispersion for neutrinos but not for photons [29, 30], and this is consistent with results from models based on causal sets [31]. Moreover, within a popular effective-theory approach to quantum-spacetime properties one describes in-vacuo dispersion for neutrinos in a way that is completely independent from that applicable to photons [32]. Given the typical energies of the astrophysical neutrinos observed by current telescopes, even just a single observation of a GRB neutrino would allow us to constrain in-vacuo dispersion for neutrinos rather tightly. However, as mentioned at the beginning, no such observation has been reported yet by studies assuming no in-vacuo dispersion.
@@
As observed in Refs. [21, 22, 23, 24, 20, 33, 34] the search of GRB neutrinos affected by in-vacuo dispersion must necessarily rely on a statistical approach: we could never be sure that a certain neutrino is a GRB neutrino, but we might at some point observe enough GRB-neutrino candidates (neutrinos that could come from a GRB, assuming in-vacuo dispersion) to be sure that at least some of them are actually GRB neutrinos affected by in-vacuo dispersion. This is due to the size of the time window that one must adopt in order to investigate the hypothesis of in-vacuo dispersion. Standard searches of GRB neutrinos (assuming $\eta=0$, no in-vacuo dispersion) can assume that the neutrino and the electromagnetic signal from the GRB travel essentially at the same speed, and therefore their GRB-neutrino candidates are looked for within a tiny time window, rendering background issues insignificant. For $\eta\neq 0$, also taking into account that the neutrino energy will have some uncertainty (and however one will inevitably always end up testing an hypothesis of $\eta$ taking values in a certain range), the difference in observation time between the neutrino and the electromagnetic signal would be sizably uncertain, and the time window for searches of GRB-neutrino candidates would have to be correspondingly large, resulting in background issues such that one could not be sure that a specific GRB-neutrino association is correct.
@@
Evidently the effectiveness of such a statistical approach depends rather crucially on the criteria used for the selection of candidate GRB neutrinos. And our main objective here is to test a novel proposal for these criteria. Testing alternative strategies of analysis on presently-available IceCube data is also important because it can set the stage for a later, more mature, phase of the research program, eventually using also data from KM3Net [35] and IceCube-Gen2 [36].
@@
Previous searches [21, 22, 23, 24, 20, 33, 34] of GRB neutrinos affected by in-vacuo dispersion looked for neutrinos within a time window of fixed size (neutrino-energy-independent) after the GRB onset, and put on the same footing both GRBs whose redshift has not been measured (then having to estimate that redshift crudely, since the conjectured effect is redshift dependent) and GRBs of known redshift. Including GRBs whose redshift has not been measured has the advantage of a larger number of candidate GRB neutrinos, but of course renders the analysis vulnerable to the assumptions made to roughly estimate the redshifts. Moreover, the fixed-size time window, while easily handled computationally, imposes a restriction of the analysis to a corresponding limited range of neutrino energies: since the sought effect grows linearly with energy, a time window adapted to energies of, say, 100 TeV will inevitably be too small for neutrinos with energy much greater than 100 TeV, and in an appropriate sense it would also be too large a time window for neutrinos of energy smaller than 100 TeV (when the time window is much wider than the one really needed by the sought effect, the analysis ends up being dominated by the background). In spite of these limitations, Ref. [34], the latest such study (of which some of us were authors), provided the estimate $\eta=21.7\pm 4.5$ with a $p$-value [^2] As will be clearer for our readers when, later in this manuscript, we derive an analogous $p$-value for our novel selection criteria, this $p$-value, which in Ref. [34] was called false alarm probability, is the probability to find at least as many (and as good) GRB-neutrino candidates as were actually found in the data under the null hypothesis of no in-vacuo dispersion ($\eta=0$), whereas $\eta=21.7\pm 4.5$ is an estimate of $\eta$ obtained from the actually found GRB-neutrino candidates. $P_{[34]}=0.007$.
@@
We here take the results of Ref. [34] as a starting point for testing a novel strategy of analysis. Rather than adopting a fixed-size time window and correspondingly restricting the neutrino-energy range, we consider all the neutrinos in the sample, regardless of the energy, but restrict the GRB catalogue to GRBs of known redshift. Any GRB and neutrino times of arrival are then regarded as compatible if they satisfy Eq. (1) for some $\eta\in\left[12.7,30.7\right]$ (the two-standard-deviation interval obtained from the estimate $\eta=21.7\pm 4.5$ found in Ref. [34]) within two standard uncertainties in the neutrino energy. As here shown in Sec. IV, the fact that, by relying exclusively on GRBs of known redshift, we find fewer GRB-neutrino candidates is more than compensated, for what concerns the overall statistical significance, by the sharper setup of the analysis which is available when the redshift of the GRBs is known.
@@ II IceCube neutrinos and GRBs of known redshift
Our analysis relies on the latest HESE data release by IceCube (7 September 2023), available at [39]. We chose to work with the HESE sample because it is the IceCube sample least contaminated by atmospheric neutrinos, especially above 50-60 TeV [40]. As done in previous analogous studies [21, 22, 23, 24, 20, 34], we restrict to shower (or cascade) events, since track events have very poor energy estimates (and of course the quality of the energy information is crucial for in-vacuo-dispersion studies) [41, 42]. An additional advantage of working with shower events is that they are even more robust against atmospheric contamination (for the shower morphology, the atmospheric neutrino flux above 1 TeV is suppressed by about an order of magnitude with respect to the track morphology [43]).
A sizable part of the effort we devoted to this study was aimed at compiling a reliable catalogue of GRBs with actually measured (and not just estimated or loosely constrained) redshifts. In fact, we found that all already available GRB catalogues either imposed further constraints on the GRB properties (narrower observation window, restriction to specific GRB observatories, GRB type, etc.) or were not sufficiently accurate (misreporting or omitting direction or redshift data, not taking into account the latest relevant GCNs, etc.). The resulting list is reported here in Appendix A. In about 90% of cases it agrees with an analogous general-purpose catalogue, not restricted to known-redshift GRBs, maintained by the IceCube collaboration [44]. The previous study [34] relied on this resource, but, after checking one-by-one each of its known-redshift entries, we noticed that some of the information relevant for our analysis was inaccurately reported, probably due to the automated population of the database. In most cases this was quickly corrected according to the web catalogue [45] maintained by Jochen Greiner at the MPE, which we found to be the most accurate and comprehensive among publicly available lists. Checking by hand all the relevant GCNs and consulting additional catalogues of known-redshift GRBs, we were then able to correct the few errors left and also identify a few known-redshift GRBs whose redshift is not reported in Greiner’s table (see Appendix A for details).
@@ III Simulated data for the statistical-significance analysis and background estimate
As done in the previous neutrino studies [^3] The interested reader can find applications of this method of statistical analysis to studies not involving neutrinos, e.g., in Refs. [38, 46, 47, 48, 49, 50]. of Refs. [21, 22, 23], we do this by producing simulated data through transformations of the real data which reliably preserve their morphology while washing away any possible correlations of the type predicted by Eq. (1) between the time of arrival, the neutrino energy, and $D(z)$.
@@
$\bullet$ the directions of known-redshift GRBs are independent of galactic longitude, but are correlated to some extent with galactic latitude, as the dust lying around the galactic plane makes it much more difficult to determine the redshift of a GRB in that region; also, we do not act on the GRB observation times because we find that the detection rate of known-redshift GRBs is declining with time, probably another selection effect due to older events having been more likely followed by the long-term, host-galaxy observations which are often needed to determine the redshift of a GRB.
As a first application of our simulated data we can estimate the overall background that is expected with the criteria adopted by our strategy of analysis. For this purpose, we generated $10^{5}$ instances of simulated data [^4] In the following, whenever we refer to simulated data, it is understood that we did all computations based on $10^{5}$ instances. and we used them to estimate the expected number of HESE shower neutrinos that in the absence of in-vacuo dispersion could accidentally be associated with a GRB of known redshift according to the time-energy and direction compatibility conditions specified above (with $\eta\in[12.7,30.7]$ and $\mathcal{S}_{cut}=1/4\pi$). We find that, on average, our criteria should pick up 1.02 such accidental GRB-neutrino associations.
@@ IV Main analysis
Equipped with the preparatory work reported in the previous sections, we are now ready to discuss our main analysis. We begin by looking for GRB-neutrino candidates, using the data described in Sec. II. For each GRB-neutrino pair selected by our directional criteria ($\mathcal{S}\geq 1/4\pi$) we describe the $\Delta t$ of Eq. (1) as the difference between the neutrino observation time [^5] If there was no in-vacuo dispersion, a GRB neutrino would be observed (nearly-)simultaneously with the GRB photons. We attribute the whole of the time-of-arrival difference to the $\Delta t$ of the neutrino, since photons observed from GRBs are of much lower energies than our neutrinos and the effect we are studying depends linearly on energy. Some quantum-spacetime scenarios predict in-vacuo dispersion with different magnitude for photons and neutrinos (and if the magnitude was much stronger for photons this assumption of our analysis would not be satisfied); it is however noteworthy that effective-field-theory approaches to quantum properties of spacetime predict that the type of pure in-vacuo dispersion here studied is actually forbidden for photons whereas it is allowed for neutrinos [17, 51]. and the observation time of the GRB, which we label as $\Delta t_{candidate}$, and then we use Eq. (1) and the known redshift of the GRB, which we label as $z_{GRB}$, to convert the requirement that $\eta$ should be within the interval $[12.7,30.7]$ into a corresponding range of allowed values for the neutrino energy. We then consider the GRB-neutrino pair a GRB-neutrino candidate if the energy of the neutrino is compatible with that energy range within two standard deviations (assuming 10% uncertainty in the energy of the neutrino [41]). We find that there are four such GRB-neutrino candidates. Fig. 1 provides a visual characterization of the fact that the properties of these four GRB-neutrino candidates match rather well the expectations of the in-vacuo-dispersion scenario of Eq. (1).
@@
The fact that our findings reported in Fig. 1 fit very well with the estimate $\eta=21.7\pm 4.5$, which resulted from the analysis of Ref. [34], is noteworthy since, because of the different criteria adopted for the selection of GRB-neutrino candidates, the two analyses only share one GRB-neutrino candidate, the one shown as a violet square in Fig. 1. The analysis leading to the estimate $\eta=21.7\pm 4.5$, in Ref. [34] ended up focusing on seven GRB-neutrino candidates, but only one of them involved a GRB with measured redshift.